Hillier Electron Microscopes Inc. issued $85 million in eight-year bonds to finance the expansion of its electron microscope manufacturing facility in Brantford, Ontario. The bonds pay 4% interest semi-annually and were issued at 87.439 to yield 6%.

Required

a. Complete the following table and prepare the related journal entries for the first two interest payments.

Date Interest Payment Interest Expense Amortization of Bond Discount Balance of Bond Discount Carrying Value of Bonds
Issuance
Payment One
Payment Two
Answer #1

Answer

--Requirement [a]

>Amounts are written on $ whole numbers (not in $ millions)

Date Interest Payment Interest Expense Amortization of Bond Discount Balance of Bond Discount Carrying Value of Bonds
[A = $ 85 millions x 4% x 6/12] [B = Last 'E' x 6% x 6/12] [C = B - A] [D = D - C] [E = E + C]
Issuance $10,676,850 $74,323,150 [$85 millions x87.439/100
Payment One $1,700,000 $2,229,695 $529,695 $10,147,156 $74,852,845
Payment Two $1,700,000 $2,245,585 $545,585 $9,601,570 $75,398,430